# Semantic Data Augmentation Enhanced Invariant Risk Minimization for Medical Image Domain Generalization

### v1

@@ Abstract
Deep learning has achieved remarkable success in medical image classification. However, its clinical application is often hindered by data heterogeneity caused by variations in scanner vendors, imaging protocols, and operators. Approaches such as invariant risk minimization (IRM) aim to address this challenge of out-of-distribution generalization. For instance, VIRM improves upon IRM by tackling the issue of insufficient feature support overlap, demonstrating promising potential. Nonetheless, these methods face limitations in medical imaging due to the scarcity of annotated data and the inefficiency of augmentation strategies. To address these issues, we propose a novel domain-oriented direction selector to replace the random augmentation strategy used in VIRM. Our method leverages inter-domain covariance as a guider for augmentation direction, guiding data augmentation towards the target domain. This approach effectively reduces domain discrepancies and enhances generalization performance. Experiments on a multi-center diabetic retinopathy dataset demonstrate that our method outperforms state-of-the-art approaches, particularly under limited data conditions and significant domain heterogeneity. The code for our method has been made publicly available.
Index Terms:
Domain Generalization, Invariant Risk Minimization, Semantic Data Augmentation, Medical Image Classification
@@ I Introduction
Deep learning has shown remarkable capabilities and broad applicability in medical image analysis [1, 2]. In the past decade, significant progress has been made in tasks such as organ segmentation, tumor classification, and disease progression prediction [3]. Despite these advancements, real-world clinical applications often require robust performance across diverse datasets from multiple institutions [4]. This emphasizes the importance of addressing data heterogeneity, which arises from variations in scanner vendors, imaging protocols, and operators [5]. Such heterogeneity leads to substantial variations in probability distributions across clinical datasets, significantly impacting the performance of deep learning models [6, 7, 8]. As a result, when deep neural network models are applied to new, unseen data from different domains, their limited generalization ability restricts their practical applicability in clinical settings, where generalization is crucial [9, 10].
@@
Domain generalization has recently attracted significant attention as a potential solution to this challenge [11, 12]. Existing approaches can be broadly categorized into data-centric, feature-centric, and model-centric methods. Data-centric methods focus on data augmentation to enhance domain diversity, including techniques such as Mixup [13] and style augmentation-based generation [14], which improve model robustness by simulating unseen scenarios [15]. Feature-centric methods, such as invariant risk minimization (IRM) [16], aim to learn domain-invariant features to improve generalization across different domains [17, 18, 19]. However, while vicinal invariant risk minimization (VIRM) [20] shows promise in large-scale datasets, its random direction selection strategy often results in inefficient or meaningless augmentations (as shown in fig. 1) in the context of medical imaging, where annotated data is scarce and expensive. Random direction selection may fail in medical image analysis, as augmentations without explicit alignment toward the target domain fail to bridge domain gaps.
@@
This issue is primarily due to the fact that medical image datasets are usually smaller in scale compared to natural images and exhibit high variability. Although other domain generalization algorithms tailored for medical image analysis, such as GDRNet [10], have been proposed, their main limitation lies in their strong specialization. This specialization makes them less adaptable to the diverse imaging modalities encountered in medical images, thereby restricting their ability to generalize across different types of medical imaging data.
@@
Figure 2: Overview of the proposed method framework. Our method includes an inter-domain covariance-guided direction selector and a distribution estimator that enhances the range of distribution. $\mathbf{x}^{i}$ denotes the input image from domain $i$, $\mathbf{z}^{i}$ denotes the feature representation of $\mathbf{x}^{i}$, $\mathbf{d}^{i}_{j}$ denotes the augment direction for domain $i$ to domian $j$, and $\tilde{z}^{i}_{j}$ denotes the augmented feature representation of $\mathbf{z}^{i}$ in domain $j$.
@@
To address these challenges, we propose a novel domain direction selector, guided by inter-domain covariance, to replace the random augmentation strategy used in VIRM. Unlike VIRM, which aims to improve the overlap of representations between domains through random augmentations, our method dynamically guides each augmentation toward the target domain. This ensures that augmented samples contribute meaningfully to reducing domain discrepancies and enhancing domain generalization. The proposed direction selector is designed to adjust the augmentation direction based on the inter-domain covariance, ensuring that the augmented samples effectively reduce domain discrepancies. The overall framework of our proposed method is illustrated in fig. 2. We evaluate our method on a medical image domain generalization dataset: Diabetic Retinopathy. Experimental results demonstrate that our approach consistently outperforms state-of-the-art methods, particularly in scenarios with limited data and significant domain heterogeneity. These findings highlight the robustness and practicality of our framework, providing a solid foundation for its broader application in medical imaging tasks.
@@ II Related Work
Domain generalization (DG) has emerged as a critical area of research with the goal of developing models that generalize well across different domains, particularly in the context of medical image datasets. A variety of techniques have been proposed in the literature to address domain shifts, which arise due to factors such as differences in imaging equipment, acquisition protocols, and patient demographics. These approaches can be broadly classified into three categories: data-centric methods, feature-centric methods, and model-centric methods. In this section, we review these categories in the context of medical image analysis and discuss the strengths and limitations of each.
@@ II-A Data-Centric Methods
Data-centric approaches focus on enhancing domain diversity through data augmentation strategies, enabling the model to learn from a broader range of scenarios. These methods simulate unseen domain variations, which helps improve generalization performance. Popular techniques include: Mixup: A data augmentation method that generates new samples by combining pairs of images and their corresponding labels, fostering improved model robustness [13]. StyleGAN: A generative approach that synthesizes diverse data variations by learning the distribution of image styles across different domains [21]. While these methods contribute to improved model performance by simulating a variety of scenarios, they often face challenges in medical image analysis, where annotated data is limited and expensive. The limited availability of annotated data in medical domains reduces the effectiveness of random data augmentation strategies.
@@ II-B Feature-Centric Methods
Feature-centric methods aim to learn domain-invariant features that generalize well across multiple domains. These methods focus on identifying common representations shared among domains to mitigate domain shifts and improve performance on unseen data. Invariant Risk Minimization (IRM): IRM seeks to minimize the risk of domain-specific variations by learning features that are invariant across different domains [16]. Domain-Adversarial Neural Networks (DANN): DANN employs adversarial training to align feature distributions between domains, encouraging the network to extract domain-invariant features [22]. VIRM: The recent method VIRM introduces a random augmentation strategy to enhance domain overlap and improve invariant feature learning [20]. While VIRM shows promise in large-scale datasets, its random direction selection often results in inefficient augmentations for medical imaging, where annotated data is scarce and highly variable. While feature-centric methods have demonstrated success in improving generalization across domains, they can struggle with medical imaging tasks where domain overlap is minimal or where data is highly heterogeneous. This often leads to suboptimal feature alignment.
@@ II-C Model-Centric Methods
Model-centric methods focus on designing novel model architectures or learning strategies to handle domain shifts more effectively. These methods typically incorporate domain-specific information into the learning process or employ meta-learning strategies to improve model adaptation. Meta-Learning: Meta-learning approaches such as Model-Agnostic Meta-Learning (MAML) aim to train models that can quickly adapt to new domains with minimal data [23]. Domain Alignment Networks: These models introduce domain-specific components into the architecture to align features across domains while preserving domain-specific characteristics [24]. Although model-centric methods offer flexibility and adaptability, they often require large-scale datasets for training and can be computationally expensive, which limits their applicability in medical imaging tasks where labeled data is scarce.
@@ II-D Domain Generalization in Medical Imaging
Recent research has focused on applying domain generalization techniques specifically to medical image analysis. These studies aim to address challenges such as variations in imaging modalities, acquisition protocols, and limited annotated data. Key advancements in this area include: GREEN[9] method utilizes a Graph Convolutional Network to model class dependencies in diabetic retinopathy grading, addressing label uncertainty and improving the classification results through a residual re-ranking mechanism. GDRNet [10] effectively reduces domain discrepancies and enhances out-of-distribution generalization, particularly under conditions of limited data and significant domain heterogeneity. However, its specificity limits its generalization to other medical imaging tasks with varying imaging principles and modalities.
@@ III-A Problem Definition
Deep learning models excel at linearizing features [25], enabling them to generate semantically different features by performing linear transformations on deep features while preserving the original label [26]. Zhu et al.[27] redefined semantic data augmentation by decomposing it into two components: semantic direction and semantic strength. The semantic direction specifies the features to be altered, enabling transformation into different semantics, while semantic strength controls the magnitude of the shift, as excessive augmentation can alter the label. Therefore, the medical semantic data augmentation module proposed in this paper is defined as:
@@
Vicinal Invariant Risk Minimization (VIRM) [20] is a domain generalization method aiming to minimize the expected loss across multiple training domains by learning domain-invariant features. VIRM introduces a Bayesian semantic data augmentation [27] strategy to enhance domain overlap and improve invariant feature learning. VIRM is formulated as the following optimization problem:
@@ III-B MedSDA Module
The MedSDA module consists of two components: the director (section III-C) and the estimator (section III-D) as shown in fig. 2. The director is responsible for selecting the augment direction $\mathbf{d}$, while the estimator estimates the augmentable distribution $p(\mathbf{\xi}|\mathbf{z}_{i})$. These components are jointly trained to minimize the loss function of the MedSDA module. The MedSDA module is designed to enhance the feature overlap between domains, improving the generalization performance of the model. Therefore, $\tilde{\mathbf{z}}_{i}$ is generated using eq. 1.
@@ III-C Director
[…]
@@ III-D Estimator
Another challenge with eq. 7 arises when domain-specific estimators are introduced, as they prevent feature sharing across domains, hindering the transformation of domain-invariant features. Although using shared estimators across domains can address this issue, it increases the model’s optimization complexity. Thus, our domain-shared SDA module is defined as:
@@
While eq. 8 estimates the distribution $p(\mathbf{\xi}|\mathbf{\mathbf{z}}_{i})$, additional constraints on the augmented features are essential to maintain label consistency. Specifically, consistency is constrained as:
@@ III-E Loss Function
Equation 10 consists of two components: the invariant risk and the loss function of the domain-shared SDA module $\mathcal{L}_{\phi}$. The loss function of the domain-shared SDA module is defined as:
ϕ=subscriptitalic-$\mathcal{L}_{\phi}=$ $-\frac{1}{2}\sum_{i=0}^{n}(1+\log(\boldsymbol{\sigma}^{2})-\boldsymbol{\sigma}^{2})+\frac{1}{2n}\sum_{l=1}^{n}(\mathbf{\hat{z}_{i}}-\mathbf{z}_{i})^{2}.$ (11)
For the first term $\mathcal{L}_{\text{IRM}}$ in eq. 10, there are several methods to implement IRM. However, our experiments align with the results of the VIRM framework, where VREx [28] was adopted as the specific implementation for IRM. VREx improves upon previous IRM formulations by introducing a penalty term for the risk variance, thus encouraging the model to reduce risk disparities across domains, enhancing its robustness to changes in the underlying distribution.
@@ IV Experiments
TABLE I: Comparison of the proposed method with the state-of-the-art methods on GRDBench [10] dataset. Best and second-best results are highlighted in Best and Second-Best, respectively.
@@
ERM [29] $0.676$ $0.326$ $0.137$ $0.640$ $0.201$ $0.067$ $0.755$ $0.324$ $0.099$ $0.649$ $0.535$ $0.174$ $0.680$ $0.347$ $0.119$
CABNet [30] $0.634$ $0.417$ $0.188$ $0.558$ $0.202$ $0.067$ $0.660$ $0.324$ $0.098$ $0.594$ $0.468$ $0.175$ $0.612$ $0.353$ $0.132$
Fishr [31] $0.794$ $0.547$ $0.302$ $0.774$ $0.232$ $0.149$ $0.804$ $0.355$ $0.190$ $0.726$ $0.517$ $0.272$ $0.774$ $0.413$ $0.228$
GDRNet [10] $0.681$ $0.371$ $0.277$ $0.694$ $0.403$ $0.261$ $0.708$ $0.283$ $0.246$ $0.659$ $0.272$ $0.224$ $0.685$ $0.332$ $0.252$
GREEN [9] $0.742$ $0.494$ $0.278$ $0.704$ $0.202$ $0.076$ $0.779$ $0.347$ $0.214$ $0.666$ $0.516$ $0.224$ $0.723$ $0.390$ $0.198$
IRM [16] $0.767$ $0.297$ $0.213$ $0.763$ $0.222$ $0.148$ $0.795$ $0.388$ $0.261$ $0.740$ $0.513$ $0.317$ $0.766$ $0.355$ $0.235$
VREx [28] $0.752$ $0.262$ $0.183$ $0.748$ $0.217$ $0.124$ $0.773$ $0.351$ $0.191$ $0.725$ $0.472$ $0.309$ $0.749$ $0.325$ $0.202$
VIRM [20] $0.774$ $0.356$ $0.220$ $0.801$ $0.240$ $0.190$ $0.804$ $0.378$ $0.283$ $0.728$ $0.448$ $0.276$ $0.776$ $0.356$ $0.242$

### v2

@@ Abstract
Deep learning has achieved remarkable success in medical image classification. However, its clinical application is often hindered by data heterogeneity caused by variations in scanner vendors, imaging protocols, and operators. Approaches such as invariant risk minimization (IRM) aim to address this challenge of out-of-distribution generalization. For instance, VIRM improves upon IRM by tackling the issue of insufficient feature support overlap, demonstrating promising potential. Nonetheless, these methods face limitations in medical imaging due to the scarcity of annotated data and the inefficiency of augmentation strategies. To address these issues, we propose a novel domain-oriented direction selector to replace the random augmentation strategy used in VIRM. Our method leverages inter-domain covariance to guide augmentation direction, guiding data augmentation towards the target domain. This approach effectively reduces domain discrepancies and enhances generalization performance. Experiments on a multi-center diabetic retinopathy dataset demonstrate that our method outperforms state-of-the-art approaches, particularly under limited data conditions and significant domain heterogeneity. The code for our method has been made publicly available at https://github.com/YaoyaoZhu19/MedVIRM.
@@ I Introduction
Deep learning has shown remarkable capabilities and broad applicability in medical image analysis [10601163, 10374392]. In the past decade, significant progress has been made in tasks such as organ segmentation, tumor classification, and disease progression prediction [7404017]. Despite these advancements, real-world clinical applications often require robust performance across diverse datasets from multiple institutions [multi_center]. This emphasizes the importance of addressing data heterogeneity, which arises from variations in scanner vendors, imaging protocols, and operators [9837077]. Such heterogeneity leads to substantial variations in probability distributions across clinical datasets, significantly impacting the performance of deep learning models [liu2020ms, wang2020domainmix, li2021simple]. As a result, when deep neural network models are applied to new, unseen data from different domains, their limited generalization ability restricts their practical applicability in clinical settings, where generalization is crucial [liu2020green, dgdr2023].
@@
Domain generalization has recently attracted significant attention as a potential solution to this challenge [wang2022generalizing, zhou2022domain]. Existing approaches can be broadly categorized into data-centric, feature-centric, and model-centric methods. Data-centric methods focus on data augmentation to enhance domain diversity, including techniques such as Mixup [zhang2017mixup] and style augmentation-based generation [li2023intra], which improve model robustness by simulating unseen scenarios [concutmix]. Feature-centric methods, such as invariant risk minimization (IRM) [arjovsky2019invariant], aim to learn domain-invariant features to improve generalization across different domains [chen2022pair, lin2022birm, li2022invariant]. However, while vicinal invariant risk minimization (VIRM) [zhu2024enlarging] shows promise in large-scale datasets, its random direction selection strategy often results in inefficient or meaningless augmentations (as shown in Fig.˜1) in the context of medical imaging, where annotated data is scarce and expensive. Random direction selection may fail in medical image analysis, as augmentations without explicit alignment toward the target domain fail to bridge domain gaps.
@@
This issue is primarily because medical image datasets are usually smaller in scale compared to natural images and exhibit high variability. Although other domain generalization algorithms tailored for medical image analysis, such as GDRNet [dgdr2023], have been proposed, their main limitation lies in their strong specialization. This specialization makes them less adaptable to the diverse imaging modalities encountered in medical images, thereby restricting their ability to generalize across different types of medical imaging data.
@@
Figure 2: Overview of the proposed method framework. Our method includes an inter-domain covariance-guided direction selector and a distribution estimator that augmentable the range of distribution. $\mathbf{x}^{i}$ denotes the input image from domain $i$, $\mathbf{z}^{i}$ denotes the feature representation of $\mathbf{x}^{i}$, $\mathbf{d}^{i}_{j}$ denotes the augment direction for domain $i$ to domian $j$, and $\tilde{z}^{i}_{j}$ denotes the augmented feature representation of $\mathbf{z}^{i}$ in domain $j$.
@@
To address these challenges, we propose a novel domain direction selector, guided by inter-domain covariance, to replace the random augmentation strategy used in VIRM. Unlike VIRM, which aims to improve the overlap of representations between domains through random augmentations, our method dynamically guides each augmentation toward the target domain. This ensures that augmented samples contribute meaningfully to reducing domain discrepancies and enhancing domain generalization. The proposed direction selector adjusts the augmentation direction based on the inter-domain covariance, ensuring that the augmented samples effectively reduce domain discrepancies. The overall framework of our proposed method is illustrated in Fig.˜2. We evaluate our method on a medical image domain generalization dataset: Diabetic Retinopathy. Experimental results demonstrate that our approach consistently outperforms state-of-the-art methods, particularly in scenarios with limited data and significant domain heterogeneity. These findings highlight the robustness and practicality of our framework, providing a solid foundation for its broader application in medical imaging tasks.
@@ II Related Work
Domain generalization (DG) has emerged as a critical area of research to develop models that generalize well across different domains, particularly in the context of medical image datasets. A variety of techniques have been proposed in the literature to address domain shifts, which arise due to factors such as differences in imaging equipment, acquisition protocols and patient demographics. These approaches can be broadly classified into three categories: data-centric methods, feature-centric methods, and model-centric methods. In this section, we review these categories in the context of medical image analysis and discuss the strengths and limitations of each.
@@ II-A Data-Centric Methods
Data-centric approaches focus on enhancing domain diversity through data augmentation strategies, enabling the model to learn from a broader range of scenarios. These methods simulate unseen domain variations, which helps improve generalization performance. Popular techniques include: Mixup: A data augmentation method that generates new samples by combining pairs of images and their corresponding labels, fostering improved model robustness [zhang2017mixup]. StyleGAN: A generative approach that synthesizes diverse data variations by learning the distribution of image styles across different domains [karras2019style]. While these methods contribute to improved model performance by simulating a variety of scenarios, they often face challenges in medical image analysis, where annotated data is limited and expensive. The limited availability of annotated data in medical domains reduces the effectiveness of random data augmentation strategies.
@@ II-B Feature-Centric Methods
Feature-centric methods aim to learn domain-invariant features that generalize well across multiple domains. These methods focus on identifying common representations shared among domains to mitigate domain shifts and improve performance on unseen data. Invariant Risk Minimization (IRM): IRM seeks to minimize the risk of domain-specific variations by learning features that are invariant across different domains [arjovsky2019invariant]. Domain-Adversarial Neural Networks (DANN): DANN employs adversarial training to align feature distributions between domains, encouraging the network to extract domain-invariant features [ganin2016domain]. VIRM: The recent method VIRM introduces a random augmentation strategy to enhance domain overlap and improve invariant feature learning [zhu2024enlarging]. While VIRM shows promise in large-scale datasets, its random direction selection often results in inefficient augmentations for medical imaging, where annotated data is scarce and highly variable. While feature-centric methods have demonstrated success in improving generalization across domains, they can struggle with medical imaging tasks where domain overlap is minimal or where data is highly heterogeneous. This often leads to suboptimal feature alignment.
@@ II-C Model-Centric Methods
Model-centric methods focus on designing novel model architectures or learning strategies to handle domain shifts more effectively. These methods typically incorporate domain-specific information into the learning process or employ meta-learning strategies to improve model adaptation. Meta-Learning: Meta-learning approaches such as Model-Agnostic Meta-Learning (MAML) aim to train models that can quickly adapt to new domains with minimal data [finn2017model]. Domain Alignment Networks: These models introduce domain-specific components into the architecture to align features across domains while preserving domain-specific characteristics [li2018domain]. Although model-centric methods offer flexibility and adaptability, they often require large-scale datasets for training and can be computationally expensive, which limits their applicability in medical imaging tasks where labeled data is scarce.
@@ II-D Domain Generalization in Medical Imaging
Recent research has focused on applying domain generalization techniques specifically to medical image analysis. These studies aim to address challenges such as variations in imaging modalities, acquisition protocols, and limited annotated data. Key advancements in this area include: GREEN[liu2020green] method utilizes a Graph Convolutional Network to model class dependencies in diabetic retinopathy grading, addressing label uncertainty and improving the classification results through a residual re-ranking mechanism. GDRNet [dgdr2023] effectively reduces domain discrepancies and enhances out-of-distribution generalization, particularly under conditions of limited data and significant domain heterogeneity. However, its specificity limits its generalization to other medical imaging tasks with varying imaging principles and modalities.
@@ III-A Problem Definition
Deep learning models excel at linearizing features [upchurch2017deep], enabling them to generate semantically different features by performing linear transformations on deep features while preserving the original label [ISDA]. Zhu et al.[zhu2024bayesian] redefined semantic data augmentation by decomposing it into two components: semantic direction and semantic strength. The semantic direction specifies the features to be altered, enabling transformation into different semantics, while semantic strength controls the magnitude of the shift, as excessive augmentation can alter the label. Therefore, the medical semantic data augmentation module proposed in this paper is defined as:
@@
Vicinal Invariant Risk Minimization (VIRM) [zhu2024enlarging] is a domain generalization method aiming to minimize the expected loss across multiple training domains by learning domain-invariant features. VIRM introduces a Bayesian semantic data augmentation [zhu2024bayesian] strategy to enhance domain overlap and improve invariant feature learning. VIRM is formulated as the following optimization problem:
@@ III-B MedSDA Module
The MedSDA module consists of two components: the director (Section˜III-C) and the estimator (Section˜III-D) as shown in Fig.˜2. The director is responsible for selecting the augment direction $\mathbf{d}$, while the estimator estimates the augmentable distribution $p(\mathbf{\xi}|\mathbf{z}_{i})$. These components are jointly trained to minimize the loss function of the MedSDA module. The MedSDA module is designed to enhance the feature overlap between domains, improving the generalization performance of the model. Therefore, $\tilde{\mathbf{z}}_{i}$ is generated using Equation˜1.
@@ III-C Director
We employ inter-domain covariance as the guiding principle for selecting augmentation directions. Let $x\in\mathcal{X}$ be the input features with domains $\mathcal{D}$, where $x$ has shape $(N,\text{out\_features})$ and $\text{domains}\in\mathcal{D}$ has shape $(N,)$. The covariance between the domains is computed as:
@@ III-D Estimator
Another challenge with Equation˜7 arises when domain-specific estimators are introduced, as they prevent feature sharing across domains, hindering the transformation of domain-invariant features. Although using shared estimators across domains can address this issue, it increases the model’s optimization complexity. Thus, our domain-shared SDA module is defined as:
@@
While Equation˜8 estimates the distribution $p(\mathbf{\xi}|\mathbf{\mathbf{z}}_{i})$, additional constraints on the augmented features are essential to maintain label consistency. Specifically, consistency is constrained as:
@@ III-E Loss Function
Equation˜10 consists of two components: the invariant risk and the loss function of the domain-shared SDA module $\mathcal{L}_{\phi}$. The loss function of the domain-shared SDA module is defined as:
$\mathcal{L}_{\phi}=$ $-\frac{1}{2}\sum_{i=0}^{n}(1+\log(\boldsymbol{\sigma}^{2})-\boldsymbol{\sigma}^{2})+\frac{1}{2n}\sum_{l=1}^{n}(\mathbf{\hat{z}_{i}}-\mathbf{z}_{i})^{2}.$ (11)
For the first term $\mathcal{L}_{\text{IRM}}$ in Equation˜10, there are several methods to implement IRM. However, our experiments align with the results of the VIRM framework, where VREx [krueger2021out] was adopted as the specific implementation for IRM. VREx improves upon previous IRM formulations by introducing a penalty term for the risk variance, thus encouraging the model to reduce risk disparities across domains, and enhancing its robustness to changes in the underlying distribution.
@@ IV Experiments
TABLE I: Comparison of the proposed method with the state-of-the-art methods on GRDBench [dgdr2023] dataset. Best and second-best results are highlighted in Best and Second-Best, respectively.
@@
ERM [vapnik1999nature] $0.676$ $0.326$ $0.137$ $0.640$ $0.201$ $0.067$ $0.755$ $0.324$ $0.099$ $0.649$ $0.535$ $0.174$ $0.680$ $0.347$ $0.119$
CABNet [he2020cabnet] $0.634$ $0.417$ $0.188$ $0.558$ $0.202$ $0.067$ $0.660$ $0.324$ $0.098$ $0.594$ $0.468$ $0.175$ $0.612$ $0.353$ $0.132$
Fishr [rame2021ishr] $0.794$ $0.547$ $0.302$ $0.774$ $0.232$ $0.149$ $0.804$ $0.355$ $0.190$ $0.726$ $0.517$ $0.272$ $0.774$ $0.413$ $0.228$
GDRNet [dgdr2023] $0.748$ $0.447$ $0.323$ $0.780$ $0.494$ $0.366$ $0.788$ $0.316$ $0.275$ $0.733$ $0.308$ $0.307$ $0.762$ $0.391$ $0.318$
GREEN [liu2020green] $0.742$ $0.494$ $0.278$ $0.704$ $0.202$ $0.076$ $0.779$ $0.347$ $0.214$ $0.666$ $0.516$ $0.224$ $0.723$ $0.390$ $0.198$
IRM [arjovsky2019invariant] $0.767$ $0.297$ $0.213$ $0.763$ $0.222$ $0.148$ $0.795$ $0.388$ $0.261$ $0.740$ $0.513$ $0.317$ $0.766$ $0.355$ $0.235$
VREx [krueger2021out] $0.752$ $0.262$ $0.183$ $0.748$ $0.217$ $0.124$ $0.773$ $0.351$ $0.191$ $0.725$ $0.472$ $0.309$ $0.749$ $0.325$ $0.202$
VIRM [zhu2024enlarging] $0.774$ $0.356$ $0.220$ $0.801$ $0.240$ $0.190$ $0.804$ $0.378$ $0.283$ $0.728$ $0.448$ $0.276$ $0.776$ $0.356$ $0.242$